# Spectral Convolutional Conditional Neural Processes

### v4

@@ Abstract
Neural processes (NPs) are probabilistic meta-learning models that map sets of observations to posterior predictive distributions, enabling inference at arbitrary domain points. Their capacity to handle variable-sized collections of unstructured observations, combined with simple maximum-likelihood training and uncertainty-aware predictions, makes them well-suited for modeling data over continuous domains. Since their introduction, several variants have been proposed. Early approaches typically represented observed data using finite-dimensional summary embeddings obtained through aggregation schemes such as mean pooling. However, this strategy fundamentally mismatches the infinite-dimensional nature of the generative processes that NPs aim to capture. Convolutional conditional neural processes (ConvCNPs) address this limitation by constructing infinite-dimensional functional embeddings processed through convolutional neural networks (CNNs) to enforce translation equivariance. Yet CNNs with local spatial kernels struggle to capture long-range dependencies without resorting to large kernels, which impose significant computational costs. To overcome this limitation, we propose the Spectral ConvCNP (SConvCNP), which performs global convolution in the frequency domain. Inspired by Fourier neural operators (FNOs) for learning solution operators of partial differential equations (PDEs), our approach directly parameterizes convolution kernels in the frequency domain, leveraging the relatively compact yet global Fourier representation of many natural signals. We validate the effectiveness of SConvCNP on both synthetic and real-world datasets, demonstrating how ideas from operator learning can advance the capabilities of NPs.
@@ 1 Introduction
Stochastic processes offer a mathematical framework for modeling systems that evolve with inherent randomness over continuous domains such as time and space. They underpin a wide range of scientific applications—from spatio-temporal climate dynamics to biological and physical systems—thereby motivating the development of machine learning methods that can learn from data generated by such stochastic phenomena (mathieu2021contrastive; vaughan2021convolutional; allen2025end; ashman2025gridded; dupont2021generative). Among classical approaches, Gaussian processes (GPs; rasmussen2006gaussian) provide a Bayesian framework with closed-form inference and uncertainty quantification. However, their cubic computational cost from matrix inversion and the difficulty of specifying suitable kernels—especially in high-dimensional settings—limit their scalability.
Motivated by the success of deep neural networks in large-scale function approximation, neural network-based alternatives have emerged. Neural processes (NPs; garnelo2018conditional; garnelo2018neural) exemplify this paradigm, combining ideas from GPs and deep learning within a meta-learning framework. By exposing the model to multiple realizations of an underlying stochastic process, each treated as a distinct task, NPs learn shared structure across tasks to parameterize a neural mapping that directly gives an approximation of the corresponding posterior predictive distribution (bruinsma2024convolutional). Once trained, the model enables efficient probabilistic predictions on new tasks without further training.
Since the introduction of conditional neural processes (CNPs, garnelo2018conditional) as the first class within the NPs family (jha2022neural), a wide range of extensions and variants have been proposed to enhance their effectiveness. One active line of research introduces explicit inductive biases into CNPs to encode the symmetries commonly present in scientific and physical domains (gordon2019convolutional; kawano2021group; holderrieth2021equivariant; huang2023practical; ashman2024translation; ashman2024approximately). Another major direction seeks to move beyond the mean-field factorized Gaussian predictive distributions to which CNPs are limited. A popular approach augments CNPs with stochastic latent variables, giving rise to the family of latent neural processes (LNPs; garnelo2018neural; louizos2019functional; wang2020doubly; foong2020meta; lee2020bootstrapping; volpp2021bayesian; wang2022bridge; wang2022learning; kim2022neural; jung2023bayesian; lee2023martingale; xu2023deep). Alternative approaches explore autoregressive prediction schemes (bruinsma2023autoregressive; nguyen2022transformer), Gaussian predictive distributions with non-diagonal covariances (bruinsma2021gaussian; markou2022practical), and quantile-based parameterizations of the predictive distribution (mohseni2023adaptive).
@@
In this work, we propose an alternative approach that represents functions in the frequency domain, inspired by the well-established observation that many natural processes exhibit energy concentration in low-frequency bands (field1987relations; ruderman1993statistics; wainwright1999scale). This characteristic allows for efficient approximation using only a subset of dominant spectral coefficients, enabling tractable computations while preserving the signal’s global structure. By parameterizing convolution kernels directly in the Fourier domain over a finite set of frequencies and leveraging the convolution theorem, our method attains large effective receptive fields without incurring prohibitive computational costs.
While spectral methods have been extensively studied in neural operator learning for partial differential equations (PDEs) (e.g., (li2020fourier; gupta2021multiwavelet; kovachki2021universal; kovachki2023neural; helwig2023group)), their application within NPs framework remains relatively unexplored. To bridge this gap, we propose the Spectral Convolutional Conditional Neural Process (SConvCNP)—a model that adopts Fourier neural operators (FNOs; li2020fourier) to realize global convolution while maintaining computational efficiency. Across a suite of synthetic and real-world benchmarks, SConvCNP performs competitively with state-of-the-art baselines, illustrating how ideas from neural operators can enhance the flexibility and performance of NPs.
@@ 2.1 Fourier Neural Operators
Neural operators (chen1995universal; li2020fourier; kovachki2023neural) are neural network architectures designed to learn mappings between function spaces rather than finite-dimensional vectors. Much like conventional feed-forward networks, they consist of stacked layers that alternate between operator-based transformations and pointwise nonlinearities. The common transformation choice is an integral operator with kernel $\kappa:{\mathbb{X}}\times{\mathbb{X}}\to{\mathbb{Y}}$, acting on an input function $v:{\mathbb{X}}\to{\mathbb{Y}}$ as
@@
where ${\mathbb{X}}={\mathbb{R}}^{d_{\mathbb{X}}}$ and ${\mathbb{Y}}={\mathbb{Y}}^{d_{\mathbb{Y}}}$ are Euclidean spaces with dimensions $d_{\mathbb{X}}$} and $d_{\mathbb{Y}}$}, respectively. While this work focuses on linear operators, nonlinear formulations have also been explored, including continuous formulations of softmax attention (ashman2024translation; calvello2024continuum). When the kernel is stationary ($\kappa\left(x,s\right)=\kappa\left(x-s\right)$), the operator simplifies to a convolution, $\mathcal{K}[v]=\kappa\ast v$, connecting neural operators to CNNs, where $\kappa$ is parameterized by learnable weights.
Typically, convolutional kernels are spatially local and possess a limited receptive field (luo2016understanding; Peng_2017_CVPR; wang2018non). Consequently, modeling long-range dependencies requires increasingly large kernels, which substantially inflates the parameter count. The Fourier Neural Operator (FNO, li2020fourier) overcomes this limitation by exploiting the convolution theorem (bracewell1966fourier; oppenheim1999discrete), which re-expresses convolution in the frequency domain as:
@@
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively. Rather than parameterizing the kernel $\kappa$ in the spatial domain, the FNO learns its representation directly in the Fourier domain. When the input function $v$ is approximately band-limited—i.e., its spectrum $\mathcal{F}[v](\xi)$ carries negligible energy for frequencies $|\xi|>\xi_{0}$—the high-frequency components can be truncated with minimal loss of information. This property, observed in many natural signals (field1987relations; ruderman1993statistics; wainwright1999scale), allows setting $\mathcal{F}[\kappa](\xi)=0$ outside the retained frequency band.
@@ 2.2 Neural Processes
Let $\mathfrak{P}$ denote the space of all $\mathbb{Y}$}-valued stochastic processes on $\mathbb{X}$}. Consider $\mathcal{P}\in\mathfrak{P}$, with $p(\cdot)$ denoting its corresponding finite-dimensional distribution. A task $\mathcal{D}$ is defined as a finite collection of input–output pairs sampled from a realization $f\sim\mathcal{P}$, partitioned into a context set and a query [^2] The NPs literature conventionally uses “target points” (with subscript $t$) to refer to prediction points. To prevent notational ambiguity with time indices, we instead adopt the terminology “query points.” set:
$\mathcal{D}=\left(\mathcal{D}_{c},\mathcal{D}_{q}\right)=\left(\left\{\left(x_{c,k},y_{c,k}\right)\right\}_{k\in\mathcal{I}_{c}},\;\left\{\left(x_{q,k},y_{q,k}\right)\right\}_{k\in\mathcal{I}_{q}}\right),$ (2)
@@
$y_{c,k}=f(x_{c,k})+\epsilon_{c,k},\quad y_{q,k}=f(x_{q,k})+\epsilon_{q,k},\quad\epsilon_{c,k},\epsilon_{q,k}\sim\mathcal{N}(0,\sigma_{0}^{2}).$
@@
Here, $\sigma_{0}>0$ is the noise scale, and $\mathcal{I}_{c},\mathcal{I}_{q}\subset\mathbb{N}$ are index sets corresponding to the context and query subsets, respectively. Neural processes (NPs; (garnelo2018conditional; garnelo2018neural)) constitute a class of models that employ neural networks to learn a mapping $\eta:\bigcup_{k=0}^{\infty}({\mathbb{X}}\times{\mathbb{Y}})^{k}\;\to\;\mathfrak{P}$, which takes a finite (possibly empty) context set $\mathcal{D}_{c}$ and gives a direct approximation to the posterior predictive process (bruinsma2021gaussian; bruinsma2024convolutional; ashman2024translation; ashman2024approximately). Let $q_{\mathcal{D}_{c}}(\cdot)$ denote the density of predictions $\eta[\mathcal{D}_{c}]\big((x_{q,k})_{k\in\mathcal{I}_{q}}\big)$. The NP approximation can then be expressed as
$q_{\mathcal{D}_{c}}\!\left(\left(y_{q,k}\right)_{k\in\mathcal{I}_{q}}\,\big|\,\left(x_{q,k}\right)_{k\in\mathcal{I}_{q}},\mathcal{D}_{c}\right)\approx p\!\left(\left(y_{q,k}\right)_{k\in\mathcal{I}_{q}}\,\big|\,\left(x_{q,k}\right)_{k\in\mathcal{I}_{q}},\mathcal{D}_{c}\right).$
@@
In this work, we focus on conditional NPs (CNPs, garnelo2018conditional), which restrict $q_{\mathcal{D}_{c}}$ to the family of mean-field Gaussians, i.e. $q_{\mathcal{D}_{c}}\!\big(\left(y_{q,k}\right)_{k\in\mathcal{I}_{q}}\mid\left(x_{q,k}\right)_{k\in\mathcal{I}_{q}},\mathcal{D}_{c}\big)=\prod_{k\in\mathcal{I}_{q}}q_{\mathcal{D}_{c}}\left(y_{q,k}\mid x_{q,k},\mathcal{D}_{c}\right)$. In general, CNPs parameterization of $q_{\mathcal{D}_{c}}\left(y_{q,k}\mid x_{q,k},\mathcal{D}_{c}\right)$ can be abstracted as a two-stage encoder-decoder pipeline (bruinsma2024convolutional; ashman2024translation; ashman2025gridded). The encoder $\varphi_{e}:{\mathbb{X}}\times\bigcup_{k=0}^{\infty}({\mathbb{X}}\times{\mathbb{Y}})^{k}\rightarrow{\mathbb{H}}$ maps the query location and context set into a latent representation, while the decoder $\varphi_{d}:{\mathbb{X}}\times{\mathbb{H}}\rightarrow\Theta$ transforms the query location and latent code into parameters $\theta\in\Theta$ of the predictive distribution $q_{\mathcal{D}_{c}}(y_{q,k}\mid x_{q,k},\mathcal{D}_{c})$.
@@
The vanilla CNP encoder summarizes the context set $\mathcal{D}_{c}$ using a permutation-invariant architecture (qi2017pointnet; zaheer2017deep). Each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is independently mapped by a shared feedforward network to a finite-dimensional embedding $\varepsilon_{c,k}$, which are subsequently aggregated—typically via mean pooling—into a single representation $\varepsilon_{c}$. Notably, this encoding is independent of the query location (i.e., $\varphi_{e}$ does not condition on the query). The decoder then fuses $\varepsilon_{c}$ with the query input through another feedforward network to produce the parameters of the Gaussian predictive distribution.
@@
$\eta\Big[\big\{(x_{c}+\tau,y_{c})\;\big|\;(x_{c},y_{c})\in\mathcal{D}_{c}\big\}\Big]\big((x_{q,k})_{k\in\mathcal{I}_{q}}\big)\;=\;\eta\big[\mathcal{D}_{c}\big]\big((x_{q,k}-\tau)_{k\in\mathcal{I}_{q}}\big),$
@@
for every $\tau\in{\mathbb{X}}$. They show that a broad family of such maps is expressed as $\eta[\mathcal{D}_{c}]=\varphi_{d}\big[\varphi_{e}\left[\mathcal{D}_{c}\right]\big]$, where the functional embedding is defined by
$\varphi_{e}[\mathcal{D}_{c}](\cdot)=\sum_{(x_{c},\,y_{c})\in\mathcal{D}_{c}}\phi(y_{c})\,\psi_{e}(\cdot-x_{c}),$ (3)
@@
and $\varphi_{d}:{\mathbb{H}}\to C_{b}({\mathbb{X}},{\mathbb{Y}})$ is a translation-equivariant decoder operating on a function space $\mathbb{H}$}. Here, $C_{b}({\mathbb{X}},{\mathbb{Y}})$ denotes the space of bounded continuous functions from $\mathbb{X}$} to $\mathbb{Y}$}; $\phi(y)=(1,y)$ [^3] For most applications, this representation suffices. Generally, $\phi(y)=(1,y,\dots,y^{M})$, where $M$ is the multiplicity of repeated inputs. See gordon2019convolutional and bruinsma2024convolutional for further discussion. and $\psi_{e}:{\mathbb{X}}\to{\mathbb{R}}$ is a continuous, strictly positive-definite kernel typically chosen to be a Gaussian.
@@
In implementing the ConvCNP, $\varphi_{e}[\mathcal{D}_{c}]$ is first evaluated on a uniform grid $\mathcal{G}\subset{\mathbb{X}}$ that spans the joint support of both the context and query locations. This yields a discretized representation $(\varphi_{e}[\mathcal{D}_{c}](x_{g}))_{x_{g}\in\mathcal{G}}$, which is then passed through $\varphi_{d}$ to produce $\big(\varphi_{d}\big[\varphi_{e}[\mathcal{D}_{c}]\big](x_{g})\big)_{x_{g}\in\mathcal{G}}$. Because the query inputs $(x_{q,k})_{k\in\mathcal{I}_{q}}$ may not lie exactly on the grid $\mathcal{G}$, ConvCNP employs an additional interpolation step to obtain predictions at query locations. Specifically
$(\theta_{q,k})_{k\in\mathcal{I}_{q}}=\bigg(\,\sum_{x_{g}\in\mathcal{G}}\varphi_{d}\big[\varphi_{e}[\mathcal{D}_{c}]\big](x_{g})\,\psi_{d}(x_{q,k}-x_{g})\bigg)_{k\in\mathcal{I}_{q}},$ (4)
where $\psi_{d}$ is another strictly positive kernel.
3 Spectral Convolutional Conditional Neural Process
The decoder $\varphi_{d}$ in ConvCNP is typically parameterized using standard CNNs such as U-Net (ronneberger2015u) or ResNet (he2016deep). These architectures employ discrete convolutional kernels—finite sets of learnable parameters that define localized filters operating over neighboring grid points. The kernel size, fixed a priori, determines the receptive field of each convolution (ding2022scaling) and is generally much smaller than the overall spatial extent of the input signals (romero2021ckconv; knigge2023modelling).

### v5

@@ Abstract
Neural Processes (NPs) are meta-learning models that learn to map sets of observations to approximations of the corresponding posterior predictive distributions. By accommodating variable-sized, unstructured collections of observations and enabling probabilistic predictions at arbitrary query points, NPs provide a flexible framework for modeling functions over continuous domains. Since their introduction, numerous variants have emerged; however, early formulations shared a fundamental limitation: they compressed the observed data into finite-dimensional global representations via aggregation operations such as mean pooling. This strategy induces an intrinsic mismatch with the infinite-dimensional nature of the stochastic processes that NPs intend to model. Convolutional conditional neural processes (ConvCNPs) address this limitation by constructing infinite-dimensional functional embeddings processed through convolutional neural networks (CNNs) to enforce translation equivariance. Yet CNNs with local spatial kernels struggle to capture long-range dependencies without resorting to large kernels, which impose significant computational costs. To overcome this limitation, we propose spectral ConvCNPs (SConvCNPs), which perform global convolution in the frequency domain. Inspired by Fourier neural operators (FNOs) for learning solution operators of partial differential equations (PDEs), our approach directly parameterizes convolution kernels in the frequency domain, leveraging the relatively compact yet global Fourier representation of many natural signals. We validate the effectiveness of SConvCNPs on both synthetic and real-world datasets, demonstrating how ideas from operator learning can advance the capabilities of NPs.
@@ 1 Introduction
Stochastic processes offer a mathematical framework for modeling systems that evolve with inherent randomness over continuous domains such as time and space. They underpin a wide range of scientific applications—from spatio-temporal climate dynamics to biological and physical systems—thereby motivating the development of machine learning methods that can learn from data generated by such phenomena (mathieu2021contrastive; vaughan2021convolutional; allen2025end; ashman2025gridded; dupont2021generative). Among classical approaches, Gaussian processes (GPs; rasmussen2006gaussian) provide a Bayesian framework with closed-form inference and uncertainty quantification. However, their cubic computational cost from matrix inversion and the difficulty of specifying suitable kernels—especially in high-dimensional settings—limit their scalability.
Motivated by the success of deep neural networks in large-scale function approximation, neural network-based alternatives have emerged. Neural processes (NPs; garnelo2018conditional; garnelo2018neural) exemplify this paradigm, combining ideas from GPs and deep learning within a meta-learning framework. By exposing the model to multiple realizations of an underlying stochastic process, each treated as a distinct task, NPs learn shared structures across tasks to parameterize a neural mapping that directly gives an approximation of the corresponding posterior predictive distribution (bruinsma2024convolutional). Once trained, the model enables efficient probabilistic predictions on new tasks without further training.
Since the introduction of conditional neural processes (CNPs, garnelo2018conditional) as the first class within the NPs family (jha2022neural), numerous extensions have been proposed to enhance their effectiveness. One prominent line of work focuses on incorporating explicit inductive biases into CNPs in order to better capture the symmetries that commonly arise in scientific applications (gordon2019convolutional; kawano2021group; holderrieth2021equivariant; huang2023practical; ashman2024translation; ashman2024approximately). Another major direction seeks to move beyond the mean-field factorized Gaussian predictive distributions to which CNPs are limited. A widely adopted approach augments CNPs with stochastic latent variables, giving rise to the family of latent neural processes (LNPs; garnelo2018neural; louizos2019functional; wang2020doubly; foong2020meta; lee2020bootstrapping; volpp2021bayesian; wang2022bridge; wang2022learning; kim2022neural; jung2023bayesian; lee2023martingale; xu2023deep). Complementary efforts explore explore autoregressive prediction schemes (bruinsma2023autoregressive; nguyen2022transformer), Gaussian predictive distributions with non-diagonal covariances (bruinsma2021gaussian; markou2022practical), and quantile-based parameterizations of the predictive distribution (mohseni2023adaptive).
@@
In this work, we pursue an alternative paradigm that represents functions in the frequency domain, inspired by the well-established observation that many natural processes exhibit energy concentration in low-frequency bands (field1987relations; ruderman1993statistics; wainwright1999scale). This property allows for efficient approximation using only a subset of dominant spectral coefficients, enabling tractable computations while preserving the signal’s global structure. By parameterizing convolution kernels directly in the Fourier domain over a finite set of frequencies and leveraging the convolution theorem, we can attain large effective receptive fields without incurring prohibitive computational costs.
While spectral methods have been extensively studied in neural operator learning for partial differential equations (PDEs, chen1995universal; kovachki2021universal; kovachki2023neural), their application within NPs framework remains relatively unexplored. To bridge this gap, we propose spectral convolutional conditional neural processes (SConvCNPs)—a class of models that adopt Fourier neural operators (FNOs; li2020fourier) to realize global spectral convolution while maintaining computational efficiency. Across a suite of synthetic and real-world benchmarks, SConvCNPs perform competitively with state-of-the-art baselines, illustrating how ideas from neural operators can enhance the flexibility and performance of NPs.
@@ 2.1 Fourier Neural Operators
Neural operators (chen1995universal; li2020fourier; kovachki2023neural) are neural network architectures designed to learn mappings between function spaces rather than finite-dimensional vectors. Much like conventional feed-forward networks, they consist of stacked layers that alternate between operator-based transformations and pointwise nonlinearities. The common transformation choice is an integral operator with kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{Y}$, acting on an input function $v:\mathcal{X}\to\mathcal{Y}$ as
@@
where $\mathcal{X}={\mathbb{R}}^{d_{x}}$ and $\mathcal{Y}={\mathbb{R}}^{d_{y}}$ are Euclidean spaces with dimensions $d_{x}$} and $d_{y}$}, respectively. While this work focuses on operators that are linear in $v$, nonlinear formulations have also been explored, including continuous formulations of softmax attention (ashman2024translation; calvello2024continuum). When the kernel is stationary, i.e. $\kappa\left(x,s\right)=\kappa\left(x-s\right)$, the operator simplifies to a convolution, $\mathcal{K}[v]=\kappa\ast v$, connecting neural operators to CNNs, where $\kappa$ is parameterized by learnable weights.
Typically, convolutional kernels are spatially local and possess a limited receptive field (luo2016understanding; Peng_2017_CVPR; wang2018non). Consequently, modeling long-range dependencies requires increasingly large kernels, which substantially inflates the parameter count. The Fourier Neural Operator (FNO, li2020fourier) overcomes this limitation by exploiting the convolution theorem (bracewell1966fourier; oppenheim1999discrete), which re-expresses convolution as:
@@
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively. Rather than parameterizing the kernel $\kappa$ in the physical domain, the FNO learns its representation directly in the Fourier domain. When the input function $v$ is approximately band-limited—i.e., its spectrum $\mathcal{F}[v](\xi)$ carries negligible energy for frequencies $\lVert\xi\rVert>\xi_{0}$—the high-frequency components can be truncated with minimal loss of information. This property, observed in many natural signals (field1987relations; ruderman1993statistics; wainwright1999scale), allows setting $\mathcal{F}[\kappa](\xi)=0$ outside the retained frequency band without significant information loss.
@@ 2.2 Neural Processes
Let $\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$ denote the set of probability measures over functions $f:\mathcal{X}\to\mathcal{Y}$. We assume a data-generating process $\mu\in\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, from which functions are sampled. A task $\mathcal{D}$ is defined as a finite collection of input–output pairs drawn from a noisy realization $f\sim\mu$, partitioned into a context set and a query [^2] The NPs literature conventionally uses the term “target points.” To avoid notational ambiguity with time indices, we instead use “query points.” set (bruinsma2024convolutional; ashman2024translation; ashman2024approximately):
$\mathcal{D}=\left(\mathcal{D}_{c},\mathcal{D}_{q}\right)=\left(\left\{\left(x_{c,k},y_{c,k}\right)\right\}_{k\in\mathcal{I}_{c}},\;\left\{\left(x_{q,k},y_{q,k}\right)\right\}_{k\in\mathcal{I}_{q}}\right),\vskip-7.11317pt$ (2)
@@
$y_{c,k}=f(x_{c,k})+\epsilon_{c,k},\quad y_{q,k}=f(x_{q,k})+\epsilon_{q,k},\quad\epsilon_{c,k},\epsilon_{q,k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{0}^{2}).$
@@
Here, $\sigma_{0}>0$ denotes the observation noise scale, and $\mathcal{I}_{c},\mathcal{I}_{q}\subset\mathbb{N}$ index the context and query sets, respectively. Neural processes (NPs; (garnelo2018conditional; garnelo2018neural)) constitute a class of models that employ neural networks to learn a mapping $\eta:\bigcup_{k=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{k}\to\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, which takes a finite (possibly empty) context set $\mathcal{D}_{c}$ and gives a stochastic process intended to approximate the Bayesian posterior induced by $\mu$. This approximation is constructed implicitly through its finite-dimensional marginals(garnelo2018neural; bruinsma2021gaussian; bruinsma2024convolutional; mathieu2023geometric). Denote ${\bm{x}}_{q}=(x_{q,k})_{k\in\mathcal{I}_{q}}$ and ${\bm{y}}_{q}=(y_{q,k})_{k\in\mathcal{I}_{q}}$. The NP approximation can then be expressed as
$q\!\left({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c}\right)\approx p\!\left({\bm{y}}_{q}\mid{\bm{x}}_{q},,\mathcal{D}_{c}\right),$
where $q({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c})$ and $p({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c})$ are the finite-dimensional distributions—evaluated at ${\bm{x}}_{q}$—of the learned process $\eta[\mathcal{D}_{c}]$ and of the true posterior process induced by $\mu$, respectively.
@@
In this work, we focus on conditional NPs (CNPs, garnelo2018conditional), which restrict $q$ to the family of mean-field Gaussians, i.e. $q\big({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c}\big)=\prod_{k\in\mathcal{I}_{q}}q\!\left(y_{q,k}\mid x_{q,k},\mathcal{D}_{c}\right)$. In general, CNPs parameterization of $q\!\left(y_{q,k}\mid x_{q,k},\mathcal{D}_{c}\right)$ can be abstracted as a two-stage encoder-decoder pipeline (bruinsma2024convolutional; ashman2024translation; ashman2025gridded). The encoder $\varphi_{e}:\mathcal{X}\times\bigcup_{k=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{k}\rightarrow\mathcal{H}$ maps $x_{q,k}$ and $\mathcal{D}_{c}$ into a latent representation, while the decoder $\varphi_{d}:\mathcal{X}\times\mathcal{H}\rightarrow\Theta$ transforms $x_{q,k}$ and latent code into parameters $\theta\in\Theta$ of the predictive distribution $q_{\theta}(y_{q,k}\mid x_{q,k},\mathcal{D}_{c})$.
@@
The vanilla CNP encoder summarizes the context set $\mathcal{D}_{c}$ using a permutation-invariant architecture (qi2017pointnet; zaheer2017deep). Each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is independently mapped by a shared feedforward network to a finite-dimensional embedding $\varepsilon_{c,k}$, which are subsequently aggregated—typically via mean pooling—into a single representation $\varepsilon_{c}$. Notably, this encoding is independent of query locations (i.e., $\varphi_{e}$ does not condition on ${\bm{x}}_{q}$). The decoder then fuses $\varepsilon_{c}$ with $x_{q,k}$ through another feedforward network to parameterize the Gaussian predictive distribution.
@@
$\eta\Big[\big\{(x_{c}+\tau,y_{c})\;\big|\;(x_{c},y_{c})\in\mathcal{D}_{c}\big\}\Big]\big((x_{q,k})_{k\in\mathcal{I}_{q}}\big)\;=\;\eta\big[\mathcal{D}_{c}\big]\big((x_{q,k}-\tau)_{k\in\mathcal{I}_{q}}\big),\vskip-2.84544pt$
@@
for every $\tau\in\mathcal{X}$. They show that a broad family of such maps is expressed as $\eta[\mathcal{D}_{c}]=\varphi_{d}\big[\varphi_{e}\left[\mathcal{D}_{c}\right]\big]$, where the functional embedding is defined by
$\varphi_{e}[\mathcal{D}_{c}](\cdot)=\sum_{(x_{c},\,y_{c})\in\mathcal{D}_{c}}\phi(y_{c})\,\psi_{e}(\cdot-x_{c}),\vskip-2.84544pt$ (3)
@@
and $\varphi_{d}:\mathcal{H}\to C_{b}(\mathcal{X},\mathcal{Y})$ is a translation-equivariant decoder operating on a function space $\mathcal{H}$. Here, $C_{b}(\mathcal{X},\mathcal{Y})$ denotes the space of bounded continuous functions from $\mathcal{X}$ to $\mathcal{Y}$; $\phi(y)=(1,y)$ [^3] For most applications, this representation suffices. Generally, $\phi(y)=(1,y,\dots,y^{M})$, where $M$ is the multiplicity of repeated inputs. See gordon2019convolutional and bruinsma2024convolutional for further discussion. and $\psi_{e}:\mathcal{X}\to{\mathbb{R}}$ is a continuous, strictly positive-definite kernel typically chosen to be a Gaussian.
@@
In implementing ConvCNPs, $\varphi_{e}[\mathcal{D}_{c}]$ is first evaluated on a uniform grid $\mathcal{G}\subset\mathcal{X}$ that spans the joint support of both ${\bm{x}}_{q}$ and ${\bm{x}}_{c}$. This yields a discretized representation $(\varphi_{e}[\mathcal{D}_{c}](x_{g}))_{x_{g}\in\mathcal{G}}$, which is then passed through $\varphi_{d}$ to produce $\big(\varphi_{d}\big[\varphi_{e}[\mathcal{D}_{c}]\big](x_{g})\big)_{x_{g}\in\mathcal{G}}$. Because ${\bm{x}}_{q}$ may not lie exactly on $\mathcal{G}$, ConvCNPs employ an additional interpolation step to obtain smooth predictions at ${\bm{x}}_{q}$. Specifically
$(\theta_{q,k})_{k\in\mathcal{I}_{q}}=\bigg(\,\sum_{x_{g}\in\mathcal{G}}\varphi_{d}\big[\varphi_{e}[\mathcal{D}_{c}]\big](x_{g})\,\psi_{d}(x_{q,k}-x_{g})\bigg)_{k\in\mathcal{I}_{q}},\vskip-2.84544pt$ (4)
where $\psi_{d}$ is another strictly positive kernel. Note that this step can be viewed as part of the decoder itself, thereby preserving the overall encoder–decoder abstraction discussed earlier.
3 Spectral Convolutional Conditional Neural Processes